### v8

@@ Abstract
Data from national accounts show no effect of change in net saving or consumption, in ratio to market-value capital, on change in growth rate of market-value capital (capital acceleration). Thus it appears that capital growth and acceleration arrive without help from net saving or consumption restraint. We explore ways in which this is possible, and discuss implications for economic teaching and public policy.
@@ 1 Introduction and overview
The data are most easily explained by the present value principle, and by production efficiencies enabled through innovation. Value is created in the mind of the market at the moment when prospective cash flows are discounted. It is created only if the market sees a path, step by step, from the start, to practical realization of those prospective cash flows. Then capital growth arrives when the market first evaluates prospective cash flows, and is realized eventually in physical outcomes insofar as the market has predicted correctly. That is to say that capital growth, in the mind of the market, precedes and pays for any needed investment in new plant and equipment.
@@ 2 Net saving, capital growth and capital acceleration
The argument has continued from growth in Eq. (2) to acceleration in Eqs. (3) and (4) on the reasoning that causal relationships among variables, in this case between thrift and capital growth rate, tend to be revealed most reliably in their concurrent changes. [^3] The ratio $\frac{\Delta A}{\Delta B}$, that is, compares changes in A to concurrent changes in B, and so suggests a causal relationship between A and B insofar as that ratio holds constant. This is the principle on which the concepts of correlation, covariance and regression are founded. Eq. (4) was derived from Eq. (3) so that success of predictions from growth assumptions can be measured against a standard of unity. For notational convenience, then, we define the "thrift index" $\theta$ by $\theta=\frac{\Delta s^{*}}{\Delta g(K)}$, and restate Eq. (4) as
@@
Test results for the predictions $\frac{s^{*}}{g(K)}=\frac{\Delta s^{*}}{\Delta g(K)}=1$, under thrift assumptions, appear in Fig. 1 and in Tables 1 and 2. Results were found as GDP-weighted averages over all countries and years, from data for net saving and market-value capital taken from national accounts. They show $\frac{s^{*}}{g(K)}=0.481$, with regression of $s^{*}$ on $g(K)$ found at 0.0771, and $\frac{\Delta s^{*}}{\Delta g(K)}=0.064$, with regression of $\Delta s^{*}$ on $\Delta g(K)$ at 0.0559.
@@
Regression of $s^{*}$ on $g(K)$ 0.0771∗∗∗
Observations 1,826
R2 0.84511
Within R2 0.10669
Year fixed effects $\checkmark$
Country fixed effects $\checkmark$
Regression of $\Delta s^{*}$ on $\Delta g(K)$ 0.0559∗∗∗
Observations 1,574
R2 0.37071
Within R2 0.27233
Year fixed effects $\checkmark$
Country fixed effects $\checkmark$
@@
Note: Thrift theory predicts $\frac{s^{*}}{g(K)}=1$. Free growth theory makes no prediction for these data.
Table 3: Average $\theta_{s}^{*}$ in 92 countries (screen = 0.01). Number of years clearing screen shown in ()
@@
Note: Thrift theory predicts $\theta_{s}^{*}=1$.
@@ 3 Interpretation of test results
$\theta_{s}$ or $\frac{\Delta s^{*}}{\Delta g(K)}$, measured at $\theta_{s}^{*}=0.064$, gives a separate measure of causal relatedness between $s^{*}$ and $g(K)$ as explained. Regression of $\Delta s^{*}$ on $\Delta g(K)$, found at 0.0559, gives a measure of causal relatedness between $\Delta s^{*}$ and $\Delta g(K)$. Might these small but positive findings allow the possibility that some capital growth is explained by net saving?
Any explanation must account for the chance that replacement saving will sometimes not be enough to offset depreciation. This shortfall is likeliest in prolonged downturns, where loss of income motivates households to forego net saving, and then to invade capital by putting off maintenance expense and replacement investment in order to protect consumption. Effects will include a positive effect on regression of $s^{*}$ on $g(K)$, and an upward effect on $\frac{\Delta s^{*}}{\Delta g(K)}$, as $s^{*}$ and $g(K)$ decline together in response to downward pressure on income. This interpretation could fit the appearance in Fig. 2 that $\frac{\Delta s^{*}}{\Delta g(K)}$, shown as $\theta^{*}_{s}$, found higher values during the period of the dot com and subprime crises.
In general, it might be expected that households stabilize consumption by building up reserves through net saving when earnings are higher, and deplete them when earnings are lower, as described in the life cycle theory of saving (Modigliani (1954)) or the permanent income hypothesis (Friedman (1957)). Free growth theory explains the rise in earnings as caused by a rise in productivity of current plant and implements and skills, and so predicts the same effects on regressions and $\frac{\Delta s^{*}}{\Delta g(K)}$ as does thrift theory, but from an opposite causality.
@@ 9 Net output
The concept of net output is found in our discussions, but not in our equations. Net output means value added. Appendix B will show an argument that to equate value added to the sum of consumption and capital growth, and neither less nor more, would miss two components, one negative and one positive, in value added to human capital, and that both of these components are unmeasurable by known means. Since we work here from measurable data for consumption and net saving and market value capital, we did our best to reason from those three quantities alone.
@@ 12 Displays
The web appendix includes displays of $\theta_{s}^{*}$ for each year in each country over the report period. These tend to show upward and downward spikes in values of $\theta_{s}^{*}$ in some years. Those spikes tend to be associated with small absolute values of denominators, in these cases $\Delta g(K)$, in those countries and years. Small denominators magnify errors in measurements of numerators. Worse, when denominators are small, small mismeasurements in them might reverse the term in sign.
To maximize reliability of test results, we apply a range of screens to omit years where absolute denominators fall below a given threshold. Some displays show data for all years, regardless of denominator size. Others screen out all years where absolute denominators are less then 0.01, then 0.025, then continuing upward in increments of 0.025 to a maximum screen of 0.15. $\theta_{s}^{*}$ is plotted for each country unscreened and at each of the seven successive levels of screening. All displays applied a screen of 0.01. The denominator whose absolute value is screened is capital acceleration $\Delta g(K)$ or capital growth rate $g(K)$ in all displays.
@@ 13 Disclaimers
We accept that capital growth is impossible without capital replacement first, to make up losses to depreciation, and that the only source of capital replacement is replacement investment. Thus we accept the necessity and efficacy of replacement investment. We dispute only the efficacy of net investment in realizing capital growth after replacement investment has made up for depreciation.
@@ Appendix A.   Equivalence of saving and investment
Any usage which treats saving and investment as equal must deal with the fact that saving held in cash is not investment in the usual sense of spending on productive capital, and contributes nothing to output. Keynes (1936) defined that unspent saving as intended saving, and actual saving as the part so spent.
We suggest that investors seek to maximize return, within risk tolerance, and will sometimes hold saving in cash or equivalents at zero return in recessions and depressions when positive returns cannot be found, and so when investing in the usual sense would tend to reduce capital and output rather than increase them. Thus investment could have the usual meaning in usual times, when prospective returns bring animal spirits, and could include saving in cash when not. It is in this sense that we equate saving to investment.

### v9

@@ Abstract
Data from national accounts show no effect of change in net saving, in ratio to market-value capital, on change in growth rate of market-value capital (capital acceleration). Thus it appears that capital growth and acceleration arrive without help from net saving. We explore ways in which this is possible, and discuss implications for economic teaching and public policy.
@@ 1 Introduction and overview
The data are most easily explained by the present value principle, by production efficiencies enabled through innovation, and by investment in the new technology from depreciation of the technology replaced. Value is created in the mind of the market at the moment when prospective cash flows are discounted. It is created only if the market sees a path, step by step, from the start, to practical realization of those prospective cash flows. Then capital growth arrives when the market first evaluates prospective cash flows, and is realized eventually in physical outcomes insofar as the market has predicted correctly. That is to say that capital growth, in the mind of the market, precedes and pays for any needed investment in new plant and equipment. [^2] Present value is time-discounted, and affords innovators only enough, at first, for first steps in implementation. As implementation proceeds, present value rises and affords more.
@@ 2 Net saving, capital growth and capital acceleration
The argument has continued from growth in Eq. (2) to acceleration in Eqs. (3) and (4) on the reasoning that causal relationships among variables, in this case between thrift and capital growth rate, tend to be revealed most reliably in their concurrent changes. Eq. (4) was derived from Eq. (3) so that success of predictions from growth assumptions can be measured against a standard of unity. For notational convenience, then, we define the "thrift index" $\theta$ by $\theta=\frac{\Delta s^{*}}{\Delta g(K)}$, and restate Eq. (4) as
@@
Test results for the predictions $\frac{s^{*}}{g(K)}=\theta=1$, under thrift assumptions, appear in Fig. 1 and in Tables 1, 2 and 3. Screens to which they refer are explained in Section 13 below. Results were found as GDP-weighted averages over all countries and years, from data for net saving and market-value capital taken from national accounts. They show $\frac{s^{*}}{g(K)}=0.481$, with regression of $s^{*}$ on $g(K)$ found at 0.0771, and $\frac{\Delta s^{*}}{\Delta g(K)}=0.064$, with regression of $\Delta s^{*}$ on $\Delta g(K)$ at 0.0559.
  Regression of $s^{*}$ on $g(K)$ 0.0771∗∗∗ (0.0144) Observations 1,826 R2 0.84511 Within R2 0.10669 Year fixed effects $\checkmark$ Country fixed effects $\checkmark$ Regression of $\Delta s^{*}$ on $\Delta g(K)$ 0.0559∗∗∗ (0.0057) Observations 1,574 R2 0.37071 Within R2 0.27233 Year fixed effects $\checkmark$ Country fixed effects $\checkmark$
@@
Note: $s^{*}$ is defined as $\frac{S_{net}}{K}$. Thrift theory predicts $\frac{s^{*}}{g(K)}=1$. Free growth theory makes no prediction for these data.
Table 3: Average $\theta$ in 92 countries (screen = 0.01). Number of years clearing screen shown in ()
@@
Note: $\theta$ is defined as $\frac{\Delta s^{*}}{\Delta g(K)}$. Thrift theory predicts $\theta=1$.
@@ 3 Interpretation of test results
$\theta$ or $\frac{\Delta s^{*}}{\Delta g(K)}$, measured at $0.064$, gives a separate measure of causal relatedness between $s^{*}$ and $g(K)$ as explained. Regression of $\Delta s^{*}$ on $\Delta g(K)$, found at 0.0559, gives a measure of causal relatedness between $\Delta s^{*}$ and $\Delta g(K)$. Might these small but positive findings allow the possibility that some capital growth is explained by net saving?
Any explanation must account for the chance that replacement saving will sometimes not be enough to offset depreciation. This shortfall is likeliest in prolonged downturns, where loss of income motivates households to forego net saving, and then to invade capital by putting off maintenance expense and replacement investment in order to protect consumption. Effects will include a positive effect on regression of $s^{*}$ on $g(K)$, and an upward effect on $\frac{\Delta s^{*}}{\Delta g(K)}$, as $s^{*}$ and $g(K)$ decline together in response to downward pressure on income. This interpretation could fit the appearance in Fig. 2 that $\theta$ found higher values during the period of the Iraq-Kuwait, dot com and subprime crises.
In general, it might be expected that households stabilize consumption by building up reserves through net saving when earnings are higher, and deplete them when earnings are lower, as described in the life cycle theory of saving (Modigliani (1954)) or the permanent income hypothesis (Friedman (1957)). Free growth theory explains the rise in earnings as caused by a rise in productivity of current plant and implements and skills, and so predicts the same effects on regressions and $\theta$ as does thrift theory, but from an opposite causality.
@@
4 Inferences for the creative destruction idea
Creative destruction (Schumpeter (1912) [^3] Published by Schumpeter in 1912, and Translated by Opie in 1934.) is obsolescence of existing capital through innovation of new and preferred technology, as with horse-drawn carriages and automobiles. Granting the obsolescence of old ways by new, it does not follow that innovation reduces values of assets already present. Data cited in this paper appear to show that capital growth is financed wholly by depreciation investment. A natural source of that investment is depreciation of assets the new technology will replace, and a natural market for the new technology may be owners of the assets to be replaced. Carriage owners were better able to afford cars when the carriages in which they had invested before lived out their economic lives as originally expected, so that the depreciation component in the cash flows originally expected can pay for cars in replacement just as it might have paid for new carriages.
This analysis suggests that market forces favor a kind through orderly obsolescence through replacement without premature displacement as the new technology awaits its turn. It also shows that the new technology as a separate entity grows by net investment financed from disinvestment from the technology replaced. It is only at the scale of the closed economy, meaning of all assets together, that net saving and investment appear to be unproductive.
@@ 10 Net output
The concept of net output is found in our discussions, but not in our equations. Net output means value added. Appendix B will show an argument that to interpret value added as the sum of consumption and capital growth, and neither less nor more, would miss two components, one negative (human depreciation) and one positive (self-invested work), in value added to human capital, and that both of these components are unmeasurable by known means. Since we work here from measurable data for net saving and market value capital, we did our best to reason from those two quantities alone.
@@ 13 Displays
The web appendix includes displays of $\theta$ for each year in each country over the report period. These tend to show upward and downward spikes in values of $\theta$ in some years. Those spikes tend to be associated with small absolute values of denominators, in these cases $\Delta g(K)$, in those countries and years. Small denominators magnify errors in measurements of numerators. Worse, when denominators are small, small mismeasurements in them might reverse the term in sign.
To maximize reliability of test results, we apply a range of screens to omit years where absolute denominators fall below a given threshold. Some displays in the web appendix show data for all years, regardless of denominator size. Others screen out all years where absolute denominators are less then 0.01, then 0.025, then continuing upward in increments of 0.025 to a maximum screen of 0.15. $\theta$ is plotted for each country unscreened and at each of the seven successive levels of screening. All displays in the main body of this paper apply a screen of 0.01. The denominator whose absolute value is screened is capital acceleration $\Delta g(K)$ or capital growth rate $g(K)$ in all displays.
@@ 14 Disclaimers
We accept that capital growth is impossible without capital replacement first, to make up losses to depreciation, and that the expected source of capital replacement is replacement investment. Thus we accept the necessity and efficacy of replacement investment. We dispute only the efficacy of net investment in realizing capital growth after replacement investment has made up for depreciation.
@@ Appendix A.   Equivalence of saving and investment
Any usage which treats saving and investment as equal must deal with the fact that saving held in cash is not investment in the usual sense of spending on new productive capital, and contributes nothing to output. Keynes (1936) defined that unspent saving as intended saving, and actual saving as the part so spent.
We suggest that investors seek to maximize return, within risk tolerance, and will sometimes hold saving in cash or equivalents at zero return in recessions and depressions when positive returns cannot be found, and so when investing in the usual sense would tend to reduce capital and output rather than increase them. Thus investment could have the usual meaning in usual times, when prospective returns bring animal spirits, and could include saving in cash when not. It is in this sense, differing from Keynes by equating intended (cash) investment to actual investment when firms and households see nothing better, that we equate saving to investment.